# Inside the black box: Neural network-based real-time prediction of US recessions

### v2

@@ 3.1.2 Long short-term memory
FFN is restricted to one-way signal flow, meaning that there is no feedback mechanism where the output of a layer can influence the same layer. Consequently, FFN lacks the ability to capture temporal dependencies in time series data. Although time series data can be fed into a FFN by incorporating additional input units representing previous time points, the main limitation lies in the fixed dimensionality of inputs and outputs. In other words, the precise length of temporal dependence must be predetermined, which is often unknown in real-world scenarios. This is where RNN comes into play. RNN establishes connections that form cycles, allowing for feedback loops where data can be fed back into the input before being forwarded again. This feedback loop enables RNN to maintain an internal state or memory to process sequences of inputs or time steps. Theoretically, RNN can retain all information over time and handle long-term dependencies. However, they face two computational challenges. Firstly, as input sequences grow longer, the backpropagation process relies heavily on the chain rule, which may lead to vanishing gradients. If any gradient approaches zero, all other gradients will diminish exponentially fast due to the multiplicative nature of the chain rule. This phenomenon, known as the vanishing gradient problem, prevents effective training in the model. Secondly, depending on the length of input sequences, the gradient of the loss function can become excessively large and result in numerical instability, referred to as the exploding gradient problem.
@@
The figure depicts a basic LSTM neural network architecture, comprising a LSTM layer and an output layer. The data is represented as a three-dimensional tensor, with dimensions for batch size, variables, and time steps. Each LSTM cell in the LSTM layer handles information retrieval at a specific time point using specialized processing gates in the hidden units.
Hochreiter and Schmidhuber (1997) introduced the long short-term memory (LSTM) architecture and a corresponding learning algorithm to address the challenges of error back-flow and long-term dependency in RNN. Figure 2 illustrates a basic LSTM neural network comprising a LSTM layer and an output layer. The input data is represented by a three-dimensional tensor, with dimensions for batch size, variables, and time steps. The LSTM layer consists of as many LSTM cells as there are time steps, with each cell responsible for information retrieval at a specific time point in a time series. These cells contain hidden units comprising special nodes and gates that are designed to process the information. Figure 3 provides a closer look at the hidden units of LSTM and GRU. For both Figure 2 and 3, the graphics of the hidden units are adapted from Chris Colah’s blog [^3] https://colah.github.io/posts/2015-08-Understanding-LSTMs/. The LSTM unit on the left variables three gates (forget, input, and output gates) denoted by red dotted lines, along with a cell state, which is a crucial distinction between RNN and LSTM network. The cell state functions like a conveyor belt that extends across the entire sequence, facilitating the retention of information over long periods. This mechanism closely resembles the long-term memory function of the human brain. The three gates regulate the flow of information to and from the cell state, enhancing the overall functionality of the network.
Figure 3: Hidden units of a LSTM and a GRU network with specialized gates
The figure displays the hidden units of LSTM and GRU network in comparison to each other. The LSTM unit has three gates (forget, input, and output) and a cell state, distinguishing it from RNN. The cell state acts as a conveyor belt for retaining information over time, similar to human long-term memory, while the gates control information flow for improved network functionality. The GRU simplifies the LSTM structure by merging two gates into a single gate and combining the cell and hidden state. It replaces the output gate with an update gate that processes new information in relation to the previous hidden state and updates it accordingly.
Upon the encounter of information from the previous time step $h_{t-1}$ with new information $x_{t}$, the forget gate evaluates which information from the cell state $C_{t-1}$ should be discarded. Formally, this can be described as follows:
@@
This equation presents a concise representation of the mathematical operations within the forget gate. The terms enclosed in brackets represent the linear operations within the activation function, introducing nonlinearity. The input vector $[h_{t-1}\enspace x_{t}]$ is multiplied by the weight matrix $W_{f}$ and combined with the bias vector $b_{f}$, which are then passed through the activation function, typically a sigmoid function. This ensures that the output values fall within the range from 0 to 1. Later, these values will be multiplied by each corresponding element in the previous cell state $C_{t-1}$, determining the proportion of old information to be retained in the new cell state.
The subsequent step involves identifying the portion of new information deemed valuable for storage in the cell state. This process occurs in two stages. Firstly, a node utilizing the hyperbolic tangent function proposes a vector of new candidate values for the cell state, denoted as $\tilde{C}_{t}$:
@@
The cell state undergoes the hyperbolic tangent function and is subsequently multiplied by the output of the output gate, yielding the new hidden state $h_{t}$. Therefore, $h_{t}$ selectively captures the most pertinent information from the cell state, representing the short-term memory, while the cell state retains the long-term memory.
@@ 3.1.3 Gated recurrent unit
In Figure 3, the right panel depicts a gated recurrent unit (GRU) which differs from the LSTM in terms of the structure and quantity of its gates. GRU, introduced by Cho et al. (2014), aims to simplify the LSTM by combining two gates into a single gate and merging the cell state with the hidden state. Unlike LSTM, GRU eliminates the output gate and integrates the functions of the forget and input gates into a single gate known as the update gate. This update gate processes new information based on the previous hidden state, consequently updating the hidden state.
@@ 3.2 Model specifications
Before estimating the model, it is necessary to prepare the chosen type of neural network. Figure 4 illustrates a simplified workflow chart outlining the process of forecasting using neural networks. Recession forecasting using neural networks involves five steps: Data must first undergo preprocessing before being fed into the model. Depending on the specific type of neural networks the network architecture is chosen. Hyperparameters are then optimized using the time series cross-validation in expanding window scheme. The second and third steps can be combined, treating the number of layers and units within those layers as hyperparameters. Once the data is prepared and the optimal set of hyperparameters is determined, the model is ready to be estimated and generate predictions.
@@
The first step involves preprocessing the data before it can be fed into the model. This preprocessing stage comprises multiple smaller steps. Initially, each series is evaluated and transformed individually to ensure stationarity. Zhang et al. (2020) argue that while advanced neural network-based optimization algorithms can handle non-stationary data, they incorporate historical gradients in the update calculations, which can result in a lack of relevance to past information when dealing with non-stationary data where distributions change over time. Therefore, it is prudent to use stationary time series data. The second step involves seasonally adjusting and standardizing all explanatory variables. Lastly, a generator function is employed to generate data batches throughout the estimation process. Data batches are sets of data that are processed all at once. Optimization algorithms based on stochastic gradient descent often utilize smaller batch sizes or minibatches. Bottou (2010) and Ge et al. (2015) highlight that minibatches offer the advantage of avoiding memory loss and introducing sufficient noise into each gradient update, aiding in escaping saddle points or local minima while achieving faster convergence.
Occasionally, data may have time delays or missing values. In such cases, a $k$-nearest-neighbor imputation method is used during the data pre-processing stage. Experimenting with alternative methods, such as the bagged tree imputation algorithm, does not affect the results significantly, although the latter is known for its superior performance on account of higher computational costs according to Stekhoven and Bühlmann (2011).
Additionally, the LSTM and GRU networks require a three-dimensional tensor as input, consisting of the batch size, the number of variables, and the number of time steps as axes. To accommodate LSTM and GRU network, the generator function transforms the data into a suitable three-dimensional tensor format. To fully utilize the capacity of neural networks, all 12 monthly lags of the predictors are included. De Veaux and Ungar (1994) argue that due to the overparameterization of neural networks, individual weights associated with multicollinearity become less influential. The final output of a neural network is a result of various combinations of activation functions that involve interactions among predictors, making the impact of multicollinearity typically insignificant. Moreover, the backpropagation algorithm used in neural networks does not require inverting matrices, which can be problematic when there is perfect or severe multicollinearity.
@@
In Section 3.1, three neural network models are presented, each based on a different architecture. The number of units in the input layer is determined by the number of variables. Depending on the desired number of time points to consider, the input dimension increases accordingly. For example, if the last 12 months of data from 25 predictors are to be used for future predictions, the input dimension would increase from 25 to $25\times 12=300$. Given the relatively small dataset with a maximum of 657 observations across 25 explanatory variables, a maximum of two hidden layers for FFN and one hidden layer for LSTM and GRU is allowed. The number of hidden layers and the number of units in those layers are treated as hyperparameters that should be optimized during the cross-validation process, which will be discussed in more detail later. Both the input and hidden layers of FFN use the rectified linear unit as the activation function for their units. The output layer consists of a single unit that uses the sigmoid activation function, producing a number between 0 and 1. This number can be interpreted as the probability of a recession.
As illustrated in Figure 2, the LSTM neural network architecture consists of as many LSTM cells as the number of time steps to look back. In the context of recession forecasting, I choose to examine the temporal variations of the predictors over the last 12 months, which corresponds to 12 lagged variables of the same predictor. Hence, there are 12 LSTM cells, each with a same number of hidden units ranging from 16 to 64. The first unit in the first LSTM cell is linked to the first unit in the second LSTM cell, and the same holds for the rest of the LSTM cells. This approach ensures that each chain of units explores different variable dimensions at different time points. Additionally, the network may require a second chain of cells, leading to a hidden or stacked LSTM layer. The optimal number of stacked layers is determined by the hyperparameter optimization process. The output layer in LSTM is the same as in FFN. GRU follows the same architecture as LSTM but differs in the structure and functionality of the hidden units within the GRU network.
@@
Deep neural networks with multiple hidden layers have the capacity to learn complex relationships. However, when training data are limited, some of these relationships may be the outcome of sampling noise and might not be present in real test data. This gives rise to overfitting, which occurs when a model fails to generalize from observed data to new data. Overfitting is a commonly acknowledged challenge in the supervised machine learning framework, and various methods employing different strategies have been devised to mitigate its effects and prevent the model from overfitting. Ying (2019) categorizes methods to address overfitting into four groups: network reduction, data augmentation, early stopping, and regularization.
@@
Complex models with numerous parameters require a substantial amount of data to distinguish meaningful patterns from noise. Therefore, expanding the training data is an effective approach to enhance the model’s generalizability. Data augmentation involves not only acquiring more training data but also employing techniques directly applied to the existing data, such as random sampling or reshuffling. I am unable to address this issue due to two reasons. Firstly, I use the same data as in chapter 2 to compare the performance of neural networks with traditional statistical models like logistic regression. Therefore, data augmentation is not suitable as it would distort the actual performance. Secondly, the temporal dependence of the time series data across multiple variables is a crucial aspect that needs to be explored. In this context, data augmentation strategies such as random sampling or reshuffling are not desirable.
@@
Instead, early stopping is implemented as a strategy to halt the training process when the gap between training loss and validation loss begins to widen. The parameters are then set to the values corresponding to the smallest gap, preventing the model from excessively memorizing the training data. This helps in improving the model’s performance on unseen data. To incorporate early stopping, a call is included during the estimation process that stops the iteration and restores the best parameters when the validation loss does not decrease for 5 consecutive epochs during cross-validation and for 10 consecutive epochs during the final estimation, where an epoch means one complete pass of the training dataset through the algorithm.
@@
Regularization is a technique used to mitigate the impact of less important variables in a model. There are two commonly applied categories of regularization methods in neural networks. The first category involves adding a penalty term or regularizer to the loss function. Two well-known types of regularizers, namely $L1$ and $L2$, are commonly used, similar to their application in penalized regression. $L1$ regularization, known as Lasso regression (Tibshirani (1996)), assigns zero weights to unimportant variables, effectively removing them from the model. This ensures that only influential variables with significant effects on the variable of interest are retained. On the other hand, $L2$ regularization, known as Ridge regression (Hoerl and Kennard (1970)), assigns lower weights to unimportant variables rather than discarding them completely. This approach aims to extract as much relevant information as possible while controlling model complexity and reducing overfitting. In my model, I incorporate $L2$ regularization to address these concerns, as chapter 2 demonstrated the superiority of Ridge regression over Lasso regression in the context of recession forecasting. The second category of regularization methods involves using a technique called dropout. Dropout was introduced by Srivastava et al. (2014) and involves randomly dropping units and their connections from the neural network during training. A certain percentage of hidden units are randomly dropped to create a thinned network, which is then trained using stochastic gradient descent. After training, the dropped units are restored, and the process is repeated. Dropout can be seen as approximating the effects of averaging the estimates over multiple smaller networks, while also preventing overfitting. I leverage the dropout technique in my model, treating the dropout percentage as a hyperparameter to be optimized through cross-validation.
@@
Apart from the hyperparameters related to the overfitting issue, neural networks also have other hyperparameters that need to be determined prior to training. These hyperparameters include the number of layers, the number of units in these layers, batch size, and learning rate. The values of these hyperparameters must be specified before training can begin. There are various methods to search for the best combination of hyperparameters, known as hyperparameter optimization. One simple approach is manual search, where the researchers select their own set of hyperparameters based on existing literature and evaluate the model performance on validation data. This process is repeated with different hyperparameter settings until the best combination is found. However, manual search can be time-consuming and does not guarantee finding the optimal hyperparameters. Grid search is another approach, where a predetermined set of values is specified for each hyperparameter. Every possible combination of values is then evaluated, resulting in a large number of trials. However, as indicated by Bellman (1961), the number of possible combinations grows exponentially with the number of hyperparameters, leading to the curse of dimensionality. This can make grid search impractical for models with many hyperparameters. A more effective approach, demonstrated by Bergstra and Bengio (2012), is random search. Random search involves randomly selecting values from predetermined ranges for each hyperparameter. The advantage of random search is that it can identify a good or even better set of hyperparameters within a smaller fraction of computation time compared to manual or grid search methods. Following this approach, the random search is conducted using predefined ranges of values for the hyperparameters. The types of hyperparameters and their specific ranges used for random search in the optimization process are reported in Table 3.
@@
The table provides the ranges of possible values for different neural network-based hyperparameters of the neural networks used in the analysis. The optimal combination of these hyperparameters is obtained by random search using the time series cross-validation technique in expanding window scheme.
Figure 5 illustrates the form of cross-validation that is employed to determine the optimal set of hyperparameters. Each block of the training set is divided into two sections at every iteration, with the validation set always coming after the training split. This approach ensures that the natural order of observations is maintained. For each block, the size of the validation set remains the same, while the training set gets larger as the validation set of the prior block is added to it. The number of iterations in the cross-validation setup depends on the size of each block. If the block size is small, the number of blocks and iterations increases. However, selecting an excessively small block size may result in certain blocks having insufficient or no recession data, leading to sampling bias. Considering the information presented in Table 2, which indicates the longest period between two recessions in US history prior to the Great Recession as 128 months, the length of each validation block is set to be 128 months to ensure that each validation set contains data from at least one recession. The length $l$ of the first training block is equal to $l=N_{total}-128\times(\left\lfloor\frac{n_{total}}{128}\right\rfloor-1)$, where $N_{total}$ denotes the total number of observations in a data vintage. This guarantees that the first training block is always larger than the validation block. Subsequently, an expanding window approach is adopted, increasing the length of the training set by 128 months at each step while maintaining a validation block of the same length (128 months). This methodology ensures that scarce recession data are fully utilized.
@@
For each combination of hyperparameters, the average validation loss across all cross-validations is computed and stored. The set of hyperparameters with the lowest average validation loss is selected as the optimal choice. According to mathematical calculations by Bergstra and Bengio (2012), using random search with 60 trials is likely to yield a set of hyperparameters that falls within the top 5% interval around the optimal solution with a 95% probability. Therefore, 60 trials are conducted for each training process. Once the optimal combination of hyperparameters is determined, the entire dataset is split into training and validation part in the same proportion, maintaining the chronological order of the data. Afterwards, the final estimation and validation of the model is performed to improve the accuracy of the parameters and check for any signs of overfitting.
@@
In most cases, there is no analytical solution available for minimizing such problems, and therefore the parameters need to be estimated numerically. To perform this estimation, I utilize the Adam optimizer, which was introduced by Kingma and Ba (2014). The Adam optimizer is specifically designed for optimizing stochastic loss functions using first-order gradients. It combines the strengths of two popular optimization methods, AdaGrad (Duchi et al. (2011)) and RMSProp (Tieleman and Hinton (2012)). The Adam optimizer offers computational efficiency and requires minimal memory resources.
@@
Since the effectiveness of predictive variables may vary across different forecast horizons, I explore five distinct windows: nowcasting, immediate-term, short-term, medium-term, and long-term. These windows correspond to predicting the current month, one month ahead, three months ahead, six months ahead, and twelve months ahead, respectively. Accounting for the two-month publication lag, the forecast horizons represent 2-, 3-, 5-, 8-, and 14-steps-ahead forecasts using the real-time set of predictive variables. For example, suppose it is November 2006, and I have data available up to September 2006, which was published in October 2006. If I aim to conduct nowcasting and forecast the recession for the current month, I use all available data up to September 2006 to predict for November 2006, resulting in a two-steps-ahead forecast. Similarly, for other forecast horizons, two extra months or steps need to be considered.
@@
After the models generate probability predictions, they are transformed into monthly zero-one indicators that represent the state of the economy, recession or boom, based on predefined cutpoints. The conventional approach is to use a fixed threshold, typically set at 0.5. Under this approach, if a probability prediction is equal to or greater than 0.5, it is classified as a recession, and vice versa. However, it is often mentioned that business cycles exhibit asymmetry, and therefore a 50% threshold may not be optimal for linear modeling methods. Berge and Jorda (2011) propose an optimal cutpoint within the range of 0.3 to 0.6 based on smoothed state probabilities estimated by Chauvet and Piger (2008). Ng (2014) suggests using different thresholds for different forecast horizons, ranging from 0.3 to 0.44. Vrontos et al. (2021) adopt a fixed threshold of 0.33 for classification and evaluation purposes. In the case of neural networks, they are assumed to be capable of capturing the asymmetry and other nonlinear characteristics of business cycles. Furthermore, they need to be compared to linear frameworks to assess their predictability. Hence, for comparison purposes, I opt to follow the traditional approach and use a fixed threshold of 0.5 for all the models in this study.
@@
The model is reestimated for the entire out-of-sample period to capture dynamic structural changes in the data. To manage computational resources, a forecast window of 12 months is selected, during which the estimated parameters remain constant. This means that once a neural network model is built, it is used to predict for the next 12 months using the latest available data. Afterward, the observed data for those 12 months is added to the existing sample, and the neural network model is reestimated using this extended dataset. The updated parameters are then used to forecast for the following 12 months, and so on. Although this repeated estimation approach is time-consuming and costly, it mimics the real-time process of generating predictions using the most up-to-date data available. To clarify this procedure, I focus on the nowcasting forecasts. The out-of-sample period begins in November 2006, and I want to predict for that month. In November 2006 data up to September 2006 are available. The training and validation set consists of 476 monthly observations from February 1967 to September 2006. Using the optimized hyperparameters obtained through time series cross-validation, I compute the nowcasting forecast for November 2006. The same model uses the most recent data available each month to generate predictions for the next 12 months. For the next iteration, data of the last 12 months are added to the training and validation set. The model is reestimated and used to predict for November 2007 and onwards. This expanding window procedure is repeated, gradually increasing the size of the training and validation set, to generate out-of-sample forecasts for the test period spanning from November 2006 to October 2021. The same approach is adopted for the other forecast horizons.

### v3

@@ 3.1.2 Long short-term memory
FFN is restricted to one-way signal flow, meaning there is no feedback mechanism where the output of a layer can influence the same layer. Consequently, FFN cannot capture temporal dependencies in time series data. Although time series data can be fed into an FFN by incorporating additional input units representing previous time points, the main limitation lies in the fixed dimensionality of inputs and outputs. In other words, the precise length of temporal dependence must be predetermined, which is often unknown in real-world scenarios. This is where RNN comes into play. RNN establishes connections that form cycles, allowing for feedback loops where data can be fed back into the input before being forwarded again. This feedback loop enables RNN to maintain an internal state or memory to process sequences of inputs or time steps. Theoretically, RNN can retain all information over time and handle long-term dependencies. However, they face two computational challenges. Firstly, as input sequences grow longer, the backpropagation process relies heavily on the chain rule, which may lead to vanishing gradients. If any gradient approaches zero, all other gradients will diminish exponentially fast due to the multiplicative nature of the chain rule. This phenomenon, known as the vanishing gradient problem, prevents effective training in the model. Secondly, depending on the length of input sequences, the gradient of the loss function can become excessively large and result in numerical instability, referred to as the exploding gradient problem.
@@
The figure depicts a basic LSTM neural network architecture comprising an LSTM layer and an output layer. The data is represented as a three-dimensional tensor, with dimensions for batch size, variables, and time steps. Each LSTM cell in the LSTM layer handles information retrieval at a specific time point using specialized processing gates in the hidden units.
Hochreiter and Schmidhuber (1997) introduced the long short-term memory (LSTM) architecture and a corresponding learning algorithm to address the challenges of error back-flow and long-term dependency in RNN. Figure 2 illustrates a basic LSTM neural network comprising an LSTM and output layer. The input data is represented by a three-dimensional tensor, with dimensions for batch size, variables, and time steps. The LSTM layer consists of as many LSTM cells as there are time steps, with each cell responsible for information retrieval at a specific time point in a time series. These cells contain hidden units comprising special nodes and gates designed to process the information. Figure 3 provides a closer look at the hidden units of LSTM and GRU. For both Figures 2 and 3, the graphics of the hidden units are adapted from Chris Colah’s blog [^3] https://colah.github.io/posts/2015-08-Understanding-LSTMs/. The LSTM unit on the left features three gates (forget, input, and output gates) denoted by red dotted lines and a cell state, a crucial distinction between RNN and LSTM networks. The cell state functions like a conveyor belt that extends across the entire sequence, facilitating the retention of information over long periods. This mechanism closely resembles the long-term memory function of the human brain. The three gates regulate the flow of information to and from the cell state, enhancing the overall functionality of the network.
Figure 3: Hidden units of an LSTM and a GRU network with specialized gates
The figure displays the hidden units of LSTM and GRU networks compared to each other. The LSTM unit has three gates (forget, input, and output) and a cell state, distinguishing it from RNN. The cell state acts as a conveyor belt for retaining information over time, similar to human long-term memory, while the gates control information flow for improved network functionality. The GRU simplifies the LSTM structure by merging two gates into a single gate and combining the cell and hidden state. It replaces the output gate with an update gate that processes new information with respect to the previous hidden state and updates it accordingly.
Upon encountering information from the previous time step $h_{t-1}$ with new information $x_{t}$, the forget gate evaluates which information from the cell state $C_{t-1}$ should be discarded. Formally, this can be described as follows:
@@
This equation presents a concise representation of the mathematical operations within the forget gate. The terms enclosed in brackets represent the linear operations within the activation function, introducing nonlinearity. The input vector $[h_{t-1}\enspace x_{t}]$ is multiplied by the weight matrix $W_{f}$ and combined with the bias vector $b_{f}$, which is then passed through the activation function, typically a sigmoid function. This ensures that the output values range from 0 to 1. Later, these values will be multiplied by each corresponding element in the previous cell state $C_{t-1}$, determining the proportion of old information retained in the new cell state.
The subsequent step involves identifying the portion of new information considered valuable for storage in the cell state. This process occurs in two stages. Firstly, a node utilizing the hyperbolic tangent function proposes a vector of new candidate values for the cell state, denoted as $\tilde{C}_{t}$:
@@
The cell state undergoes the hyperbolic tangent function and is multiplied by the output gate output, yielding the new hidden state $h_{t}$. Therefore, $h_{t}$ selectively captures the most pertinent information from the cell state, representing the short-term memory, while the cell state retains the long-term memory.
@@ 3.1.3 Gated recurrent unit
In Figure 3, the right panel depicts a gated recurrent unit (GRU), which differs from the LSTM regarding the structure and quantity of its gates. GRU, introduced by Cho et al. (2014), aims to simplify the LSTM by combining two gates into a single gate and merging the cell state with the hidden state. Unlike LSTM, GRU eliminates the output gate and integrates the forget and input gates functions into a single gate known as the update gate. This update gate processes new information based on the previous hidden state, consequently updating the hidden state.
@@ 3.2 Model specifications
Before estimating the model, it is necessary to prepare the chosen type of neural network. Figure 4 illustrates a simplified workflow chart outlining neural networks’ forecasting process. Recession forecasting using neural networks involves five steps: Data must undergo preprocessing before being fed into the model. The network architecture is chosen depending on the specific type of neural network. Hyperparameters are optimized using the time series cross-validation in the expanding window scheme. The second and third steps can be combined, treating the number of layers and units within those layers as hyperparameters. Once the data is prepared and the optimal set of hyperparameters is determined, the model can be estimated, and predictions can be generated.
@@
The first step involves preprocessing the data before it can be fed into the model. This preprocessing stage comprises multiple smaller steps. Initially, each series is evaluated and transformed individually to ensure stationarity. Zhang et al. (2020) argue that while advanced neural network-based optimization algorithms can handle non-stationary data, they incorporate historical gradients in the update calculations, which can result in a lack of relevance to past information when dealing with non-stationary data where distributions change over time. Therefore, it is prudent to use stationary time series data. The second step involves seasonally adjusting and standardizing all explanatory variables. Lastly, a generator function generates data batches throughout the estimation process. Data batches are sets of data that are processed all at once. Optimization algorithms based on stochastic gradient descent often utilize smaller batch sizes or minibatches. Bottou (2010) and Ge et al. (2015) highlight that minibatches offer the advantage of avoiding memory loss and introducing sufficient noise into each gradient update, aiding in escaping saddle points or local minima while achieving faster convergence. Additionally, the LSTM and GRU networks require a three-dimensional tensor as input, consisting of the batch size, the number of variables, and the number of time steps as axes. To accommodate the LSTM and GRU network, data is transformed into a suitable three-dimensional tensor format.
@@
Section 3.1 presents three neural network models, each based on a different architecture. The number of variables determines the number of units in the input layer. The input dimension increases accordingly depending on the desired number of time points to consider. The number of hidden layers and units in those layers are treated as hyperparameters that should be optimized during cross-validation, which will be discussed in more detail later. Both the input and hidden layers of FFN use the rectified linear unit as the activation function for their units. The output layer consists of a single unit that uses the sigmoid activation function, producing a number between 0 and 1. This number can be interpreted as the probability of a recession.
As illustrated in Figure 2, the LSTM neural network architecture consists of as many LSTM cells as the number of time steps to look back. In the context of recession forecasting, I examine the temporal variations of the predictors over the last 12 months, corresponding to 12 lagged variables of the same predictor. De Veaux and Ungar (1994) argue that due to the overparameterization of neural networks, individual weights associated with multicollinearity become less influential. The final output of a neural network results from various combinations of activation functions that involve interactions among predictors, making the impact of multicollinearity typically insignificant. Moreover, the backpropagation algorithm used in neural networks does not require inverting matrices, which can be problematic when perfect or severe multicollinearity exists. Hence, there are 12 LSTM cells, each with the same number of hidden units ranging from 16 to 64. The first unit in the first LSTM cell is linked to the first in the second LSTM cell, and the same holds for the rest of the LSTM cells. This approach ensures that each chain of units explores different variable dimensions at different time points. Additionally, the network may require a second chain of cells, leading to a hidden or stacked LSTM layer. The hyperparameter optimization process determines the optimal number of stacked layers. The output layer in LSTM is the same as in FFN. GRU follows the same architecture as LSTM but differs in the structure and functionality of the hidden units within the GRU network.
@@
Deep neural networks with multiple hidden layers have the capacity to learn complex relationships. However, when training data are limited, some of these relationships may result from sampling noise and might not be present in real test data. This gives rise to overfitting, which occurs when a model fails to generalize from observed data to new data. Overfitting is a commonly acknowledged challenge in the supervised machine learning framework, and various methods employing different strategies have been devised to mitigate its effects and prevent the model from overfitting. Ying (2019) categorizes methods to address overfitting into three groups: network reduction, early stopping, and regularization.
@@
Early stopping is implemented to halt the training process when the gap between training loss and validation loss widens. The parameters are then set to the values corresponding to the smallest gap, preventing the model from memorizing the training data excessively. This helps in improving the model’s performance on unseen data. To incorporate early stopping, a call is included during the estimation process that stops the iteration and restores the best parameters when the validation loss does not decrease for 5 consecutive epochs during cross-validation and for 10 consecutive epochs during the final estimation, where an epoch means one complete pass of the training dataset through the algorithm.
@@
Regularization is a technique to mitigate the impact of less important variables. There are two commonly applied categories of regularization methods in neural networks. The first category involves adding a penalty term or regularizer to the loss function. Two well-known types of regularizers, namely $L1$ and $L2$, are commonly used, similar to their application in penalized regression. $L1$ regularization, known as Lasso regression (Tibshirani (1996)), assigns zero weights to unimportant variables, effectively removing them from the model. This ensures that only influential variables with significant effects on the variable of interest are retained. On the other hand, $L2$ regularization, known as Ridge regression (Hoerl and Kennard (1970)), assigns lower weights to unimportant variables rather than discarding them completely. This approach aims to extract as much relevant information as possible while controlling model complexity and reducing overfitting. The second category of regularization methods involves using a technique called dropout. Dropout was introduced by Srivastava et al. (2014) and involves randomly dropping units and their connections from the neural network during training. A certain percentage of hidden units are randomly dropped to create a thinned network, which is then trained using stochastic gradient descent. After training, the dropped units are restored, and the process is repeated. Dropout approximates the effects of averaging the estimates over multiple smaller networks while preventing overfitting. I implement the dropout technique in my model, treating the dropout percentage as a hyperparameter to be optimized through cross-validation.
@@
Besides the hyperparameters related to the overfitting issue, neural networks also have other hyperparameters that must be determined prior to training. These hyperparameters include the number of layers, the number of units in these layers, batch size, and learning rate. The values of these hyperparameters must be specified before training can begin. There are various methods to search for the best combination of hyperparameters, known as hyperparameter optimization. A straightforward approach is a manual search, where the researchers select their own set of hyperparameters based on existing literature and evaluate the model performance on validation data. This process is repeated with different hyperparameter settings until the best combination is found. However, manual search can be time-consuming and does not guarantee finding the optimal hyperparameters. Grid search is another approach where a predetermined set of values is specified for each hyperparameter. Every possible combination of values is then evaluated, resulting in many trials. However, as indicated by Bellman (1961), the number of possible combinations grows exponentially with the number of hyperparameters, leading to the curse of dimensionality. This can make grid search impractical for models with many hyperparameters. A more effective approach, demonstrated by Bergstra and Bengio (2012), is random search. The random search involves randomly selecting values from predetermined ranges for each hyperparameter. The advantage of random search is that it can identify a good or even better set of hyperparameters within a smaller fraction of computation time compared to manual or grid search methods. Following this approach, the random search is conducted using predefined ranges of values for the hyperparameters. The types of hyperparameters and their specific ranges used for random search in the optimization process are reported in Table 3.
@@
The table provides the ranges of possible values for different neural network-based hyperparameters of the neural networks used in the analysis. The optimal combination of these hyperparameters is obtained by random search using the time series cross-validation technique in an expanding window scheme.
Figure 5 illustrates the form of cross-validation employed to determine the optimal set of hyperparameters. Each block of the training set is divided into two sections at every iteration, with the validation split always coming after the training split. This approach ensures that the natural order of observations is maintained. For each block, the size of the validation set remains the same, while the training set gets larger as the validation set of the prior block is added to it. The number of iterations in the cross-validation setup depends on the size of each block. If the block size is small, the number of blocks and iterations increases. However, selecting an excessively small block size may result in certain blocks having insufficient or no recession data, leading to sampling bias. Considering the information presented in Table 2, which indicates the longest period between two recessions in US history prior to the Great Recession as 128 months, the length of each validation block is set to be 128 months to ensure that each validation set contains data from at least one recession. The length $l$ of the first training block is equal to $l=N_{total}-128\times(\left\lfloor\frac{n_{total}}{128}\right\rfloor-1)$, where $N_{total}$ denotes the total number of observations in a data vintage. This guarantees that the first training block is always larger than the validation block. Subsequently, an expanding window approach is adopted, increasing the length of the training set by 128 months at each step while maintaining a validation block of the same length (128 months). This methodology ensures that scarce recession data are fully utilized.
@@
The average validation loss across all cross-validations is computed and stored for each combination of hyperparameters. The optimal choice is the set of hyperparameters with the lowest average validation loss. According to mathematical calculations by Bergstra and Bengio (2012), using a random search with 60 trials will likely yield a set of hyperparameters within the top 5% interval around the optimal solution with a 95% probability. Therefore, 60 trials are conducted for each training process. Once the optimal combination of hyperparameters is determined, the entire dataset is split into training and validation parts in the same proportion, maintaining the chronological order of the data. Afterward, the final estimation and validation of the model are performed to improve the accuracy of the parameters and check for any signs of overfitting.
@@
In most cases, no analytical solution is available for minimizing such problems, so the parameters must be estimated numerically. I utilize the Adam optimizer introduced by Kingma and Ba (2014) to perform this estimation. The Adam optimizer is designed to optimize stochastic loss functions using first-order gradients. It combines the strengths of two popular optimization methods, AdaGrad (Duchi et al. (2011)) and RMSProp (Tieleman and Hinton (2012)). The Adam optimizer offers computational efficiency and requires minimal memory resources.
@@
Since the effectiveness of predictive variables may vary across different forecast horizons, I explore five distinct windows: nowcasting, immediate-term, short-term, medium-term, and long-term. These windows correspond to predicting the current month, one month ahead, three months ahead, six months ahead, and twelve months ahead, respectively.
@@
After the models generate probability predictions, they are transformed into binary indicators that represent the state of the economy, recession, or boom based on predefined cutpoints. The conventional approach uses a fixed threshold, typically set at 0.5. Under this approach, if a probability prediction equals or exceeds 0.5, it is classified as a recession, and vice versa. However, it is often mentioned that business cycles exhibit asymmetry, and therefore, a 50% threshold may not be optimal for linear modeling methods. Berge and Jorda (2011) propose an optimal cutpoint within the range of 0.3 to 0.6 based on smoothed state probabilities estimated by Chauvet and Piger (2008). Ng (2014) suggests using different thresholds for different forecast horizons, ranging from 0.3 to 0.44. Vrontos et al. (2021) adopt a fixed threshold of 0.33 for classification and evaluation purposes. In the case of neural networks, they are assumed to be capable of capturing the asymmetry and other nonlinear characteristics of business cycles. Furthermore, they must be compared to linear frameworks to assess their predictability. Hence, for comparison purposes, I follow the traditional approach and use a fixed threshold of 0.5 for all the models in this study.
@@
The model is reestimated for the entire out-of-sample period to capture dynamic structural changes in the data. To manage computational resources, a forecast window of 12 months is selected, during which the estimated parameters remain constant. This means that once a neural network model is built, it is used to predict for the next 12 months using the latest available data. Afterward, the observed data for those 12 months is added to the existing sample, and the neural network model is reestimated using this extended dataset. The updated parameters are then used to forecast for the following 12 months. Although this repeated estimation approach is time-consuming and costly, it mimics the real-time process of generating predictions using the most up-to-date data available. This expanding window procedure is repeated, gradually increasing the size of the training and validation set to generate out-of-sample forecasts for the test period spanning from November 2006 to October 2021. The same approach is adopted for the other forecast horizons.